# Barrier relaxations of the classical and quantum optimal transport problems

### v1

@@ 4.2. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    (4.18) $-\log\det(I+Z)=\sum_{j=1}^{\infty}\frac{(-1)^{j}\operatorname{Tr}Z^{j}}{j}.$
@@
(a) Let $\lambda_{i}(Z),i\in[n]$ be the eignevalues $Z$. Clearly, $\det(I+Z)=\prod_{i=1}^{n}(1+\lambda_{i}(Z))$. If $\|Z\|_{2}<1$ then $|\lambda_{i}(Z)|<1$ for $i\in[n]$. Recall that for a complex $t,|t|<1$ one has the Taylor expansion $-\log(1+t)=\sum_{j=1}^{\infty}\frac{(-1)^{j}t^{j}}{j}$. Hence
@@
Let $\rho_{i}\in\mathrm{H}_{n_{i},++,1},i\in[2]$. Consider the minimum (1.7), which corresponds to the maximum dual problem (4.9). Apply the IPM short step interior path algorithm with the barrier (4.14) to the $-\min$ problem as in (2.2), where $r$ is given by (4.13). Choose a starting point given by (4.12). The CPA algorithm finds the value $\kappa(C,R)$ within precision $\delta>0$ in
@@ 5.1. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    The interior of the ball $\mathrm{B}((X_{1,0},\ldots,X_{d,0}),r)=\{\sum_{j=1}^{d}\|X_{j}-X_{j,0}\|_{F}^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((X_{1,0},\ldots,X_{d,0}),r)$, contains $(X_{1},,\ldots,X_{d})$ that satisfies the conditoins (5.10).
@@
The following proposition is a genralization of Proposition 4.3.
@@
$\frac{d^{d}\varepsilon}{2}\log\left(d\|C\|_{2}^{2}+d\big{(}\frac{2\|C\|_{2}}{\prod_{i=1}^{d}\lambda_{n}(\rho_{i})}\big{)}^{2}\right)$
@@
Let $\rho_{i}\in\mathrm{H}_{n,++,1},i\in[d]$. Consider the minimum (1.7), which corresponds to the maximum dual problem (5.9). Apply the IPM short step interior path algorithm with the barrier (4.14) to the $-\min$ problem as in (2.2), where $r$ is given by (5.13). Choose a starting point given by (4.12). The CPA algorithm finds the value $\kappa(C,R)$ within precision $\delta>0$ in $O\big{(}n^{d/2}\log\frac{rn^{d}}{\delta}\big{)}$ iterations. The arithmetic complexity of CPA is $O\big{(}n^{7d/2}\log\frac{n^{d}r}{\delta}\big{)}$.
The proof of this Theorem is similar to the proof Theorem 4.5. We discuss briefly the arithmetic complexity of the IPM algorithm. Note that the matrix $C(X_{1},\ldots,X_{d})$ is $n^{d}\times n^{d}$. Hence, its inverse is computed in $O(n^{3d})$ flops. The computation of the gradient and the Hessian of $-\log(r^{2}-\sum_{i=1}^{d}\|X_{i}\|_{F}^{2})+\log r^{2}$ needs $O(dn^{2})$ and $O((dn)^{2})$ flops respectively. It is left to discuss teh computation of the gradient and Hessian $\beta$.
@@
$-\frac{\partial\log\det C(A_{1}+X_{1},\ldots A_{d}+X_{d})}{\partial x_{k_{1},i_{1}}}=\sum_{j_{2}=\ldots=j_{d}=1}^{n}e_{(i_{1},j_{2},\ldots,j_{d})(k_{1},j_{2},\ldots,j_{d})}.$
@@ Appendix A The interior point method
$\|\mathbf{x}\|\stackrel{{\scriptstyle\mathclap{\mbox{def}}}}{{=}}\|\mathbf{x}\|_{2},$
@@
A second example of a strongly standard self-concordant function for $\mathrm{S}_{,++}$ is $-\log\det X$, where

### v2

@@ 4.2. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    (4.19) $\log\det(I+Z)=\sum_{j=1}^{\infty}\frac{(-1)^{j}\operatorname{Tr}Z^{j}}{j}.$
@@
(a) Let $\lambda_{i}(Z),i\in[n]$ be the eigenvalues $Z$. Clearly, $\det(I+Z)=\prod_{i=1}^{n}(1+\lambda_{i}(Z))$. If $\|Z\|_{2}<1$ then $|\lambda_{i}(Z)|<1$ for $i\in[n]$. Recall that for a complex $t,|t|<1$ one has the Taylor expansion $\log(1+t)=\sum_{j=1}^{\infty}\frac{(-1)^{j}t^{j}}{j}$. Hence
@@
Let $\rho_{i}\in\mathrm{H}_{n_{i},++,1},i\in[2]$. Consider the minimum (1.7), which corresponds to the maximum dual problem (4.9). Apply the IPM short step interior path algorithm with the barrier (4.14) to the $\min$ problem as in (2.2), where $r$ is given by (4.13). Choose a starting point given by (4.12). The CPA algorithm finds the value $\kappa(C,R)$ within precision $\delta>0$ in
@@ 5.1. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    The interior of the ball $\mathrm{B}((X_{1,0},\ldots,X_{d,0}),r)=\{\sum_{j=1}^{d}\|X_{j}-X_{j,0}\|_{F}^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((X_{1,0},\ldots,X_{d,0}),r)$, contains all the points $(X_{1},,\ldots,X_{d})$ that satisfies the conditoins (5.10).
@@
The following proposition is a generalization of Proposition 4.3.
@@
$n^{d}\varepsilon\log\left(\frac{2\|C\|_{2}}{\prod_{i=1}^{d}\lambda_{n}(\rho_{i})}\right)$
@@
Let $\rho_{i}\in\mathrm{H}_{n,++,1},i\in[d]$. Consider the minimum (1.7), which corresponds to the maximum dual problem (5.9). Apply the IPM short step interior path algorithm with the barrier (5.14) to the $\min$ problem as in (2.2), where $r$ is given by (5.13). Choose a starting point given by (5.12). The CPA algorithm finds the value $\kappa(C,R)$ within precision $\delta>0$ in $O\big(n^{d/2}\log\frac{rn^{d}}{\delta}\big)$ iterations. The arithmetic complexity of CPA is $O\big(n^{7d/2}\log\frac{n^{d}r}{\delta}\big)$.
The proof of this Theorem is similar to the proof Theorem 4.5. We discuss briefly the arithmetic complexity of the IPM algorithm. Note that the matrix $C(X_{1},\ldots,X_{d})$ is $n^{d}\times n^{d}$. Hence, its inverse is computed in $O(n^{3d})$ flops. The computation of the gradient and the Hessian of $\log(r^{2}-\sum_{i=1}^{d}\|X_{i}\|_{F}^{2})+\log r^{2}$ needs $O(dn^{2})$ and $O((dn)^{2})$ flops respectively. It is left to discuss the computation of the gradient and Hessian $\beta$.
@@
$\frac{\partial\log\det C(A_{1}+X_{1},\ldots A_{d}+X_{d})}{\partial x_{k_{1},i_{1}}}=\sum_{j_{2}=\ldots=j_{d}=1}^{n}e_{(i_{1},j_{2},\ldots,j_{d})(k_{1},j_{2},\ldots,j_{d})}.$
@@
Acknowledgment
Part of this work was supported by IPAM in UCLA, and conducted when the author was visiting the Workshops I and III in Non-commutative Optimal Transport Program, Spring 2025.
@@ Appendix A The interior point method
A second example of a strongly standard self-concordant function for $\mathrm{S}_{,++}$ is $\log\det X$, where